OpenStudy (anonymous):

When x = 3 and y = 5, by how much does the value of 3x2 – 2y exceed the value of 2x2 – 3y ? F. 4 G. 14 H. 16 J. 20 K. 50

RhondaSommer (rhondasommer):

3x^2

RhondaSommer (rhondasommer):

or 3X2

OpenStudy (anonymous):

thanx

RhondaSommer (rhondasommer):

nonono

RhondaSommer (rhondasommer):

i am asking the question. Is the equation 3x squared?

OpenStudy (anonymous):

yes

pooja195 (pooja195):

Ok take what thy gave you and plug it into the equations 3x^2 – 2y exceed the value of 2x^2 – 3y \[\huge~\rm~\bf~3(3)^2-2(5)=?\] \[\huge~\rm~\bf~2(3)^2-3(5)=? \]

OpenStudy (anonymous):

ok

pooja195 (pooja195):

Can you solve the 2 equations?

OpenStudy (anonymous):

yes

pooja195 (pooja195):

Ok solve them and tell me what you get :)

OpenStudy (anonymous):

2^2

pooja195 (pooja195):

Not quite.. 1) 3^2=?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

(2) 2^2

OpenStudy (anonymous):

sory 3^2

pooja195 (pooja195):

Not quite...lets solve the first one =(2)(9)−(3)(5) =18−(3)(5) =18−15 18-15=?

OpenStudy (anonymous):

3

pooja195 (pooja195):

Good so 1) 3 Now can you try to solve the 2nd one? \[\huge~\rm~\bf~2(3)^2-3(5)=?\] This one.

OpenStudy (anonymous):

3^2

pooja195 (pooja195):

Are you familiar with how to solve these?

OpenStudy (anonymous):

no

pooja195 (pooja195):

Ok first lets get rid of the exponent \[\huge~\rm~\bf~3^2=3 \times 3=? \]

OpenStudy (anonymous):

then the 2

pooja195 (pooja195):

Nope 3 x 3=?

OpenStudy (anonymous):

9

pooja195 (pooja195):

Good now we have \[\huge~\rm~\bf~(2)(9)−(3)(5) \] 2 x 9=? 3 x 5-?

OpenStudy (anonymous):

18 and 15

pooja195 (pooja195):

Good :) 18-15=?

OpenStudy (anonymous):

3

pooja195 (pooja195):

Good so one of our numbers is 3 using the same method try this one \[ \huge~\rm~\bf~3(3)^2−2(5)=?\]

OpenStudy (anonymous):

3

pooja195 (pooja195):

\[\huge~\rm~\bf~3(3)^2−2(5)=17\] \[\huge~\rm~\bf~2(3)^2−3(5)=3\] Ok now do 17-3=? They want to know how much it exceeded we can find out by subtracting 17-3=?

OpenStudy (anonymous):

14

pooja195 (pooja195):

Good and that would be our final answer :)

OpenStudy (anonymous):

yes

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

thanx

pooja195 (pooja195):

yw ^_^

OpenStudy (anonymous):

;]

pooja195 (pooja195):

;-)

OpenStudy (anonymous):

^/^